### v4

@@ Appendix F Proof of Near-Optimal Regret without Knowing $C_{p,q}$
From the previous deduction, we conclude that $\mathbb{E}[s_{i}(T)]\leq l_{0}+l_{1}+\frac{5\pi^{2}}{6}$, where $l_{0}=\max\{\lfloor e^{\frac{1}{2\epsilon}}\rfloor,\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\rceil\}=O(\log T)$ and $l_{1}=\max\{l_{0},\lceil\frac{8+\epsilon}{\chi_{M}}\log T^{2}\rceil\}$. If we consider the case when replacing $C_{p,q}$ by $f(T)$ as long as T is large enough, $\chi_{M}$(an upper bound of $x$) here actually becomes a $T$-relevant value satisfying three inequalities, while the variation does not rely on the index of a certain arm since those requirements are established for all arms to attain. Now we only need to figure out the order of $\frac{1}{\chi_{M}}$ to prove our statement.
@@
This indicates we only require $c_{0}[1-c_{4}\chi_{M}^{q-1}\tau(\chi_{M})]^{\frac{1}{q}}\geq c_{1}c_{2}\chi_{M}^{1-\frac{1}{q}}+c_{1}c_{3}\chi_{M}^{1-\frac{1}{q}}\tau(\chi_{M})^{\frac{1}{q}}$ to achieve (49). Simple analysis shows that left side $\geq c_{0}[1-c_{4}\chi_{M}^{q-1}\tau(\chi_{M})]$ while the right side $\leq c_{1}c_{2}\chi_{M}^{1-\frac{1}{q}}+c_{1}c_{3}\chi_{M}^{1-\frac{1}{q}}\tau(\chi_{M})$. Thus we only require that
@@
Furthermore, by comparing (51) with (52), they simultaneously hold if
@@
Finally, through comparison between (50) and (53), we have $\frac{1}{\chi_{M}}=O(f(T)^{\max\{\frac{q}{pD},\frac{q}{p-q}\}})=O(f(T)^{\frac{q}{pD}})$. By setting $l_{0}=\max\{t_{0},\lfloor e^{\frac{1}{2\epsilon}}\rfloor,\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\rceil\}$, we have $l_{1}=O(f(T)^{\frac{q}{pD}}\log T)$ and $\mathcal{R}_{\pi}^{\mathcal{F}}(T)=O(l_{0}+l_{1}+\frac{5\pi^{2}}{6})=O(f(T)^{\frac{q}{pD}}\log T)$. $\hfill\qed$
@@ Appendix G Proof of Theorem 2
Taking the limit $\Delta_{i}\to 0$, (56) becomes $\left(6\sqrt{2}+9\right)(\epsilon+8)<912$, which holds for $\epsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$. By continuity, we conclude that (55) holds for sufficiently small $\Delta_{i}$ if $\epsilon<\frac{280-16\sqrt{2}}{2\sqrt{2}+3}$. $\hfill\qed$
@@ Appendix H Proof of Theorem 4
If $\Delta_{i}$ of arm $i$ is large, that is, if (10) and (11) directly lead to (12), then $\mathbb{E}[s_{i}(T)]/\log T$ is upper bounded by a constant not related to $\Delta_{i}$. We are interested in the case where $\Delta_{i}$ is small, i.e., when (12) becomes the main restraint of $\chi_{M}$:
@@
With the lower bound in the non-parametric setting in Burnetas and Katehakis [6] and Theorem 3, the proof is finished immediately. $\hfill\qed$
@@ Appendix I Proof of Theorem 5
Choose $L_{0}=\lceil\max\{\frac{4a_{1}\log T}{(\theta_{1}-\theta_{i})^{2}},\frac{2a_{2}\log T}{\theta_{1}-\theta_{i}}\}\rceil$, $\forall k\geq L_{0}$, we have
@@
A direct substitution of the above into (1) completes the proof. $\hfill\qed$
@@
Since $\frac{8\alpha\zeta}{x}\leq\frac{32\zeta}{x}$ as $\alpha\leq 4$ and $\frac{2\alpha x}{u_{0}(x-\zeta u_{0})}\leq\frac{16}{u_{0}}$ as $x\geq 2\zeta u_{0}$, (60) is proved. $\hfill\qed$

### v5

@@ Appendix F Proof of Near-Optimal Regret without Knowing $C_{p,q}$
From the previous deduction, we conclude that $\mathbb{E}[s_{i}(T)]\leq l_{0}+l_{1}+\frac{5\pi^{2}}{6}$, where $l_{0}=\max\{\lfloor e^{\frac{1}{2\epsilon}}\rfloor,\lceil(8+\epsilon)\log T^{2}\frac{8\log T^{2}+1}{\epsilon\log T^{2}-1}\rceil\}=O(\log T)$ and $l_{1}=\max\{l_{0},\lceil\frac{8+\epsilon}{\chi_{M}}\log T^{2}\rceil\}$. If we consider the case of replacing $C_{p,q}$ by $f(T)$ as long as $T$ is large enough, $\chi_{M}$(an upper bound of $x$) here actually becomes a $T$-relevant value satisfying three inequalities, while the variation does not depend on the index of a specific arm since those requirements are established for all arms to satisfy. Now we only need to figure out the order of $\frac{1}{\chi_{M}}$ to prove our statement.
@@
This indicates that we only require $c_{0}[1-c_{4}\chi_{M}^{q-1}\tau(\chi_{M})]^{\frac{1}{q}}\geq c_{1}c_{2}\chi_{M}^{1-\frac{1}{q}}+c_{1}c_{3}\chi_{M}^{1-\frac{1}{q}}\tau(\chi_{M})^{\frac{1}{q}}$ to achieve (49). A simple analysis shows the left side $\geq c_{0}[1-c_{4}\chi_{M}^{q-1}\tau(\chi_{M})]$ while the right side $\leq c_{1}c_{2}\chi_{M}^{1-\frac{1}{q}}+c_{1}c_{3}\chi_{M}^{1-\frac{1}{q}}\tau(\chi_{M})$. Therefore, we only require that
@@
Furthermore, by comparing (51) with (52), they will simultaneously hold if
@@
[…]
@@ Appendix G Proof of Theorem 2
[…]
@@ Appendix H Proof of Theorem 4
If $\Delta_{i}$ of arm $i$ is large, that is, if (10) and (11) directly lead to (12), then $\mathbb{E}[s_{i}(T)]/\log T$ is upper bounded by a constant not related to $\Delta_{i}$. We are interested in the case where $\Delta_{i}$ is small, i.e., when (12) becomes the main constraint of $\chi_{M}$:
@@
[…]
@@ Appendix I Proof of Theorem 5
Choose $L_{0}=\lceil\max\{\frac{4a_{1}\log T}{(\theta_{1}-\theta_{i})^{2}},\frac{2a_{2}\log T}{\theta_{1}-\theta_{i}}\}\rceil$. We have, $\forall k\geq L_{0}$,
@@
[…]
@@
[…]